# LightPAL: Lightweight Passage Retrieval for Open Domain Multi-Document Summarization

### v1

@@ Abstract
Open-Domain Multi-Document Summarization (ODMDS) is crucial for addressing diverse information needs, which aims to generate a summary as answer to user’s query, synthesizing relevant content from multiple documents in a large collection. Existing approaches that first find relevant passages and then generate a summary using a language model are inadequate for ODMDS. This is because open-ended queries often require additional context for the retrieved passages to cover the topic comprehensively, making it challenging to retrieve all relevant passages initially. While iterative retrieval methods have been explored for multi-hop question answering (MQA), they are impractical for ODMDS due to high latency from repeated large language model (LLM) inference for reasoning. To address this issue, we propose LightPAL, a lightweight passage retrieval method for ODMDS that constructs a graph representing passage relationships using an LLM during indexing and employs random walk instead of iterative reasoning and retrieval at inference time. Experiments on ODMDS benchmarks show that LightPAL outperforms baseline retrievers in summary quality while being significantly more efficient than an iterative MQA approach.
@@ 1 Introduction
Figure 1: Overview of LightPAL, a lightweight retrieval method for Open-domain Multi-Document Summarization (ODMDS). To retrieve passages required for abstract information needs, LightPAL performs the following processes: 1) Passage Graph Construction: Evaluate the relevance between passages using LLM generation probabilities and create edges between passages offline. 2) Retrieval: After retrieving initial passages using an off-the-shelf retriever, perform a Random Walk on the Passage Graph to retrieve informative context passages that are referenced by many passages. LightPAL can lightly retrieve passages that are the context for already retrieved passages without iterative reasoning and retrieval, while existing off-the-shelf retrievers cannot do so without iteration.
@@
Summarization is a task that aims to generate text that preserves the most informative content from input documents. As an advanced version of the summarization task, Open-Domain Multi-Document Summarization (ODMDS) Giorgi et al. (2023) has recently been introduced, where the inputs are a large collection of documents, and the summary should be an answer to a query that reflects the information needs of the user. For example, given a query such as “What is the story of THE RECRUIT, an SF novel?”, a summarization system needs to synthesize information from multiple sections of the novel, which may be scattered across different documents within a large collection. In this way, ODMDS requires: (1) generating a summary in response to queries that ask for an overview or open-ended questions, i.e., queries that demand a summary as an answer, and (2) synthesizing the content of multiple documents related to the query within the large document collection. Users typically have diverse information needs, making it impractical to precompile a document collection that caters to all possible queries. Consequently, the ODMDS task plays a vital role in information-seeking scenarios where the system must adapt to the specific requirements of each user.
@@
A practical approach to ODMDS is the retrieve-then-summarize framework Giorgi et al. (2023), where a retrieval model first finds relevant passages for the query, and then a language model generates a summary using the query and retrieved passages. Producing a good summary crucially depends on retrieving comprehensive and relevant information in the first step, as low coverage or irrelevant information leads to poor summaries. Existing work Zhou et al. (2023) has employed sparse and pre-trained dense retrievers, similar to open-domain question answering Zhu et al. (2021). However, these retrievers are inadequate for the ODMDS because the information needs in ODMDS are abstract since they involve overviews or open-ended questions. When the information need is abstract, the query is not expressed as a concrete question like a factoid. Consider again the query, “What is the story of THE RECRUIT, a SF novel?” (see Figure 1). This query entails potential information needs such as “Who are the characters?”, but without reading the story, it is impossible to know the entities needed to concretize such questions to search relevant passages. Therefore, existing retrievers search for passages containing the novel title “THE RECRUIT,” but they cannot search for information that addresses potential information needs such as “Who is Wayne in THE RECRUIT?” which arises after reading the passage. To answer such abstract information needs using existing retrievers, recursive search is required, where initially retrieved information is used for reasoning to search for further information.
@@
Iterative retrieval and reasoning methods have been primarily proposed for multi-hop question answering (MQA), with recent methods leveraging the reasoning capability of large language models (LLMs). For example, Trivedi et al. (2023) let LLM generate Chain-of-Thought Wei et al. (2022) to make queries to retrieve new documents, while Khalifa et al. (2023) traverse the graph of hyperlinked Wikipedia paragraphs along the path that maximized the generation probability of the given query. In this study, we applied these approaches to ODMDS for the first time and confirmed that both retrieval performance and summary quality improve in the ODMDS task (see Section 5). However, these methods are impractical for ODMDS because the number of passages relevant to the query is larger than that of MQA. This is due to the fact that ODMDS requires diverse information to answer overview queries. For instance, in the Story dataset for ODMDS benchmark Zhou et al. (2023), the average number of relevant passages is 9, while in the HotPotQA Yang et al. (2018) dataset for Multi-hop QA, at most two passages are required for an answer. Consequently, numerous iterations are necessary to gather the relevant passages for the ODMDS task, and performing LLM inference at each iteration leads to high latency. Therefore, a more efficient approach is needed to balance summary quality and computational cost for ODMDS.
@@
In this paper, we propose LightPAL, a lightweight passage retrieval method for ODMDS. It constructs a graph representing passage relationships using a LLM during indexing and then employs random walk on this graph via Personalized PageRank (PPR) Jeh and Widom (2003a) instead of iterative reasoning and retrieval at inference time. During indexing, the LLM evaluates if a passage could serve as the context for another, creating links between passages with high generation probabilities. The generation probabilities of the LLM can capture relevance between passages by considering their diverse styles and domains in the large passage collection. At retrieval, an initial set of relevant passages is retrieved using an arbitrary model. PPR then calculates the probability of reaching other passages from the initial set via a random walk on the graph. Highly probable passages are retrieved as additional context, enabling low-latency retrieval of diverse, relevant information without LLM inference during runtime.
We evaluated LightPAL on two benchmarks: ODSum Zhou et al. (2023) for ODMDS and Querysum Liu et al. (2024) for query-focused multi-document summarization (qMDS), which we adapted for ODMDS by sampling irrelevant passages from the same dataset to create a large passage collection, as qMDS assume only query-relevant passages. In the retrieve-then-summarize approach, we evaluated the quality of generated summaries in terms of relevance to reference summaries, diversity, and comprehensiveness using LLMs. Since no conventional retrieval methods specifically designed for the ODMDS task, we compared LightPAL with a pre-trained dense retriever and confirmed that our proposed method outperforms it as an unsupervised method. Additionally, we compared LightPAL to PromptRank Khalifa et al. (2023), an unsupervised iterative reranking method for MQA. Our method demonstrated favorable summary performance while achieving significantly lower latency, about 1k times faster than PromptRank  [^1] Mistral-7B-v0.1 is used as an LLM in PromptRank..
@@ 2 Related Works
2.1 Open-domain Multi-document Summarization
Open-Domain Multi-Document Summarization (ODMDSGiorgi et al. (2023); Zhou et al. (2023) is a task that involves extracting and summarizing diverse information from multiple documents into a concise and informative summary. It generalizes the query-focused summarization (QFS) and multi-document summarization (MDS) tasks. Traditionally, MDS task Ma et al. (2023), with applications like summarizing news articles Fabbri et al. (2019), scientific literature reviews Lu et al. (2020), and product reviews Wang and Ling (2016); Ganesan et al. (2010), focused on creating a single summary from a predefined set of documents. To cater to different user information needs, QFS Kulkarni et al. (2020); Liu et al. (2024), was introduced, where a user query is added to generate summaries relevant to the query.
@@ 2.1 Open-domain Multi-document Summarization
Building upon this, Giorgi et al. (2023) introduced ODMDS as a task that involves retrieving relevant documents from open-domain information sources and summarizing them to answer a given query. This research created a pseudo-ODMDS dataset using existing MDS data to evaluate the capabilities of a retrieve-then-summarize pipeline. The results revealed a decline in summary quality due to the retrieval phase. More recently, Zhou et al. (2023) created a benchmark for the ODMDS task and evaluated the performance of LLMs as summarization models. Although this work has focused on the impact of different retrieval models on summary quality, there has been no work that enables structured retrieval by associating passages at the indexing phase to address queries seeking an overview.
2.2 Retrieval for Multi-hop/document QA
Several approaches have been developed for Multi-hop Question Answering (MQA) tasks Mavi et al. (2022), which involve gathering and integrating information from multiple sources to obtain an answer. These methods typically involve iterative reasoning and retrieval. Das et al. (2019) utilizes the internal state of the reader model to update the query vector and rerank passages. Qi et al. (2019) uses the reader to extract text spans from the retrieved passages, and then uses these text spans as queries to retrieve additional passages.
Recently, retrieval using an LLM iteratively is commonly used for MQA tasks. Khalifa et al. (2023) proposes an unsupervised re-ranking method for MQA that focuses on exploring relevant hyperlinked passages and assigning them higher scores based on the probabilities of generating the given query by LLMs. Trivedi et al. (2023) utilizes the Chain-of-Thoughts technique for retrieving potentially related passages, and Wang et al. (2024) builds a passage graph from structured documents and traverses the graph to retrieve them by using an LLM agent iteratively. The iterative methods are impractical for ODMDS regarding latency, as mentioned in Section 5.3.
2.3 Retrieval for Long-form QA
Long-form Question Answering (LFQAFan et al. (2019) is similar to ODMDS, requiring paragraph-length responses to complex, open-ended questions. In LFQA, multiple pieces of evidence must be retrieved from a large corpus to answer the query. Existing research divides retrieval systems into non-interactive and interactive methods. Non-interactive methods Su et al. (2022); Krishna et al. (2021) use off-the-shelf dense/sparse retrievers to search the corpus with the initial query. Interactive methods Nakano et al. (2021); Qin et al. (2023) learn an agent that performs interactive web searches like humans. Recent web agent research by Qin et al. (2023) aims to deepen topic understanding and create informative answers by issuing follow-up questions using relevant searched information, similar to human behaviors. Our method shares this motivation but avoids their drawbacks: (1) requiring web-browsing behaviors to train the agent and (2) being inapplicable to non-web-based retrieval systems. Our unsupervised approach can be applied to any retrieval model, offering a wider range of applications.
@@ 3 ODMDS Task
Task Formulation. Open-Domain Multi-Document Summarization (ODMDS) is formally defined as follows: given a natural language query $q$ and a large collection of passages $D=\{d_{i}\}_{i}$, generate a summary $S$ corresponding to the query $q$.
Approach. Following prior work Zhou et al. (2023), we adopt retrieve-then-summarize approach for the ODMDS task  [^2] While some LLMs, such as Gemini Pro 1.5(https://deepmind.google/technologies/gemini/pro/), can handle up to 1M input tokens, the size of the passage collection is generally unbounded, making retrieval a necessity.. First, we retrieve a set of passages $D_{q}\subset D$ relevant to the query $q$ from the collection $D$. Then, we input the retrieved text and query to a language model $LM_{sum}$ to generate an abstractive summary $S=LM_{sum}(q,D_{q})$.
@@ 4 LightPAL: Proposed Method
We propose a retrieval method for addressing open-ended ODMDS queries. Our approach retrieves relevant and diverse passages while avoiding iterative retrieval and language model inference at runtime. As shown in Figure 1, The method consists of two phases:
  1. 1.
    Graph Construction: Pre-constructing a graph representing passage relationships using a large language model (LLM), linking passages with high conditional generation probabilities as potential context for a passage.
  2. 2.
    Retrieval: Retrieving an initial set of relevant passages with a conventional retriever, then employing random walk via Personalized PageRank (PPR) Jeh and Widom (2003a) on the pre-built graph. PPR calculates probabilities of reaching other passages from initial passages, retrieving highly probable ones as additional context without runtime language model inference.
4.1 Passage Graph Construction
The motivation behind our approach is that passages extracted from a larger document may not be self-explanatory and require context from other passages to be correctly understood. Even cohesive passages can benefit from related texts to deepen topic understanding. Naive retrieval systems that only consider query-passage relevance cannot capture these relationships. We aim to address this by creating an graph annotated with passage relationships using the conditional generation probabilities of LLMs. Concretely, for each ordered pair of passages $(d_{i},d_{j})$ in passage set $D$, we calculate the conditional generation probability of passage $d_{j}$ given passage $d_{i}$ as a context score using a language model:
@@ 4.1 Passage Graph Construction
4.2 Passage Retrieval
Our proposed method consists of two phases for passage retrieval: 1) Initial Passage Retrieval: Retrieve passages $D_{init}$ directly relevant to the query using any sparse/dense retrieval model. 2) Context Passage Retrieval: Using the initial passages, additionally retrieve context passages $D_{context}$ from the passage graph $G$ that provide contextual information to the initial passages. Finally, we obtain $D_{q}$ as the union of $D_{init}$ and $D_{context}$.
For context passage retrieval, we employ Personalized PageRank (PPR) Jeh and Widom (2003b). PPR calculates the probability of reaching other passages from the initial set $D_{init}$ via a random walk on the graph $G$, highly ranking passages cohesively referenced by many passages. Starting a random walk from the query-relevant initial passages, we expect to explore topics related to the query, and then retrieve highly informative context passages for those topics using PPR. This approach enables low-latency retrieval of relevant information without iterative search using LLM inference during runtime.
@@ 5.1 Settings
Data. To evaluate our approach, we used three datasets: (1) Story dataset from ODSum Zhou et al. (2023), where passages are sections from independent stories, requiring retrieval of cohesive story sections relevant to the query. (2) Meeting dataset from ODSum, where queries are questions about meeting contents. We divided each long meeting transcript (approx. 8,000 tokens) into smaller chunks of around 1,000 characters, treating each chunk as a passage to avoid irrelevant information degrading performance. If meeting transcript is labeled as relevant to query, chunks are also labeled as relevant. We used test set from both Story and Meeting. (3) Querysum Liu et al. (2024), a dataset for query-focused multi-document summarization, consists of general non-factoid questions as queries and Wikipedia paragraphs as associated passages. We selected queries with long reference summaries (70+ words) and multiple relevant passages to test complex questions. Since Querysum does not provide a passage collection $D$ for searching, we created a collection $D$ by randomly extracting 5k Wikipedia paragraphs that are not related to the selected queries and combining them with the associated passages. This ensures that the collection contains a sufficient number of irrelevant passages. To reduce the computational cost of summary generation, we truncated each passage to its first 2,000 characters, assuming that the entity information needed to answer queries is primarily found in the beginning of paragraphs. We refer to the dataset created using the above procedure as QsumOD and use it in our experiments. The statistics of the datasets are shown in Table 1.
Dataset Story Meeting QsumOD
# Queries 260 131 90
# Passages 1190 13074 5803
# PassLen. 3170 898 1391
# SummLen. 1286 1081 675
# Rel. 9.4 170.9 8.9
Rel. % 0.8% 1.3% 0.2%
Table 1: Statistics of the datasets. Abbreviations used: Rel. (Average Relevant Passages), PassLen. (Average Passage Length in characters), SummLen. (Average Reference Summary Length in characters), Rel. % (Relevant Passages Ratio).
@@
Baselines for Passage Retrieval. To investigate the impact of retrieval methods on the final summary quality, we prepared: 1) A pre-trained dense retriever that embeds passages and queries into the same vector space and retrieves the top-K passages most similar to the query. We used the bge-large-en-v1.5  [^3] https://huggingface.co/BAAI/bge-large-en-v1.5 model for embedding as it is a top-tier model in Massive Text Embedding Benchmark  [^4] https://github.com/embeddings-benchmark/mteb 2) PromptRank Khalifa et al. (2023), an unsupervised reranker for Multi-hop QA, which performs iterative reasoning and retrieval by traversing links between passages. PromptRank requires links between passages, so to enable a fair comparison, we employ the passage graph created by our method. Mistral-7B-v0.1  [^5] https://huggingface.co/mistralai/Mistral-7B-v0.1 is used as a LLM in PromptRank. For a fair comparison, both PromptRank and our proposed method initially extracted the top $|D_{init}|$ passages from the $K$ passages retrieved by the dense retriever. Each method, PromptRank and our proposed method, then retrieved the remaining $|D_{context}|=K-|D_{init}|$ context passages using their respective approaches. We set $|D_{init}|$ as top 60% of $K$.

### v2

@@ Abstract
Open-Domain Multi-Document Summarization (ODMDS) is the task of generating summaries from large document collections in response to user queries. This task is crucial for efficiently addressing diverse information needs from users. Traditional retrieve-then-summarize approaches fall short for open-ended queries in ODMDS tasks. These queries often require broader context than initially retrieved passages provide, making it challenging to retrieve all relevant information in a single search. While iterative retrieval methods has been explored for multi-hop question answering (MQA), it’s impractical for ODMDS due to high latency from repeated LLM inference. Accordingly, we propose LightPAL, a lightweight passage retrieval method for ODMDS. LightPAL leverages an LLM to pre-construct a graph representing passage relationships, then employs random walk during retrieval, avoiding iterative LLM inference. Experiments demonstrate that LightPAL outperforms naive sparse and pre-trained dense retrievers in both retrieval and summarization metrics, while achieving higher efficiency compared to iterative MQA approaches.
@@ 1 Introduction
Figure 1: LightPAL: lightweight retrieval method for Open-domain Multi-Document Summarization (ODMDS). It consists of two main processes: 1) Offline Graph Construction: evaluates passage relevance using LLM conditional generation probabilities to create edges between passages. 2) Online Retrieval: uses an off-the-shelf retriever for initial passages, then performs a Random Walk on the constructed graph to retrieve informative context passages referenced by many others. This approach efficiently retrieves context passages without iterative LLM inference in runtime.
@@
Summarization is a task that aims to generate text that preserves the most informative content from input documents. As an advanced version, Open-Domain Multi-Document Summarization (ODMDS) Giorgi et al. (2023); Zhou et al. (2023) has recently been introduced, where the inputs are a large collection of documents, and the summary should be an answer to a query that reflects the information needs of the user. ODMDS requires: (1) generating a summary in response to queries that ask for an open-ended questions, i.e., queries that demand a summary as an answer. (2) synthesizing the content of multiple documents related to the query within the large document collection. Users typically have diverse information needs, making it impractical to precompile a document collection that caters to all possible queries. Consequently, ODMDS task plays a vital role in information-seeking scenarios where the system must adapt to information need of each user.
@@
A practical approach to ODMDS is retrieve-then-summarize Giorgi et al. (2023), where a retrieval model first finds relevant passages for the query, and then a language model generates a summary. The quality of the summary heavily depends on retrieving comprehensive and relevant information. Existing work Zhou et al. (2023) has employed sparse and pre-trained dense retrievers, similar to open-domain question answering Zhu et al. (2021). However, these retrievers are often insufficient for ODMDS. This is because users seek summaries to efficiently acquire knowledge about unfamiliar topics, with information needs not typically expressed as specific queries like factoid. For instance, consider the question "Why do airlines schedule tight layovers?" (see Figure 1). Users might only realize they need to ask about "minimum connection times" after reading initial results. For such open-ended questions, recursive search is required, where initially retrieved information is used for reasoning to search for further information.
@@
Iterative retrieval and reasoning methods have been primarily proposed for multi-hop question answering (MQA), with recent methods leveraging the reasoning capability of large language models (LLMs) Trivedi et al. (2023); Khalifa et al. (2023). We applied these approaches to ODMDS, demonstrating improved retrieval performance and summary quality (Section 5). However, ODMDS presents unique challenges due to the larger number of relevant passages required for open-ended questions compared to MQA. For example, the Story dataset for ODMDS Zhou et al. (2023) averages 9 relevant passages per query, whereas HotPotQA Yang et al. (2018) for MQA typically requires only two. This necessitates numerous iterations in ODMDS, leading to high latency when using LLM inference at each iteration. Consequently, a more efficient method is needed to balance summary quality and computational cost in the ODMDS context.
@@
In this paper, we propose LightPAL, a lightweight passage retrieval method for ODMDS. It constructs a graph representing passage relationships using a LLM during indexing, then employs random walk via Personalized PageRank (PPR) Jeh and Widom (2003a) for retrieval. During indexing, the LLM evaluates if a passage could serve as context for another, creating links based on conditional generation probabilities. This captures relevance between passages by considering diverse styles and domains. At retrieval, an initial set of relevant passages is obtained using an arbitrary model. PPR then calculates the probability of reaching other passages from the initial set on the graph. This enables low-latency retrieval of diverse, relevant information without LLM inference during runtime.
We evaluated LightPAL using two benchmarks: ODSum Zhou et al. (2023) for ODMDS and LFRQA Han et al. (2024) for long-form question answering in retrieval-augmented generation (RAG). Experiment results showed that LightPAL improved retrieval performance by up to 10% compared to sparse/pre-trained dense retrievers. Results also indicated that LightPAL performed comparably to or outperformed PromptRank Khalifa et al. (2023), an unsupervised iterative reranking method for MQA, while achieving about 1,000 times lower search latency  [^1] Qwen2.5-3B is used as an LLM in PromptRank.. For generated summary quality, LightPAL outperformed baseline retrievers when retrieval performance was substantially enhanced, as measured by ROUGE metrics and GPT-4-based evaluation Liu et al. (2023). We conducted human annotations on a subset of dataset to validate the GPT-4 evaluation, which revealed a moderate alignment between GPT-4 and human assessments.
@@ 2 Related Works
Open-domain Multi-document Summarization. Open-Domain Multi-Document Summarization (ODMDSGiorgi et al. (2023); Zhou et al. (2023) is a task that summarizing information from large collection of documents into a summary as answer to given a query. Giorgi et al. (2023) introduced ODMDS and created a pseudo-ODMDS dataset using existing multi-document summarization data to evaluate the capabilities of a retrieve-then-summarize pipeline. The results revealed a decline in summary quality due to the retrieval phase. More recently, Zhou et al. (2023) created a benchmark for the ODMDS task and evaluated the performance of LLMs as summarization models. These initial studies have examined basic retrieval and summarization components of RAG systems for ODMDS. However, no research has addressed the development of iterative retrieval methods for open-ended queries in ODMDS or their efficiency.
Retrieval for Multi-hop/document QA. Multi-hop Question Answering (MQA) tasks Mavi et al. (2022) require integrating information from multiple passages to obtain an answer. Existing approaches typically employ iterative reasoning and retrieval for multi-hop searches. Das et al. (2019) update query vectors using the reader model’s state for reranking passages, while Qi et al. (2019) extract text spans from retrieved passages to query additional passages. Recent methods leverage LLMs to assess relevance between queries and passages. Khalifa et al. (2023) proposed an unsupervised re-ranking method, scoring hyperlinked passages based on LLMs’ query generation probability. As a more advanced method, researchers use LLMs as agents to traverse pre-constructed graphs of passages. Knowledge Graph Prompting Wang et al. (2024a) uses LLMs to generate query-answering information and select similar neighboring passages. GraphReader Li et al. (2024) extends this for exploration planning, relevance evaluation, and content selection. Despite effectiveness, these methods face real-world limitations due to high latency from iterative LLM inference.
Indexing for Comprehensive Question Answering in RAG Framework. Recent RAG frameworks organize and condense documents prior to retrieval for comprehensive question answering. RAPTOR Sarthi et al. (2024) constructs tree structures through recursive summarization, while GraphRAG Edge et al. (2024) creates entity graphs for hierarchical dataset summaries. These methods, like our approach, improve indexing to use contextual information in retrieval. However, they are unsuitable for our aim as they do not address the ODMDS task or target latency reduction in retrieval. GraphRAG is unsuitable due to its extensive use of LLMs for relevance scoring against multiple summaries, leading to high latency. RAPTOR, designed for single-document QA, does not apply to ODMDS scenarios where passages are from various documents.
@@ 3 ODMDS Task
Open-Domain Multi-Document Summarization (ODMDS) is formally defined as follows: given a natural language query $q$ and a large collection of passages $D=\{d_{i}\}_{i}$, generate a summary $S$ corresponding to the query $q$. Following prior work Zhou et al. (2023), we adopt retrieve-then-summarize approach for the ODMDS task  [^2] While some LLMs, such as Gemini Pro 1.5(https://deepmind.google/technologies/gemini/pro/), can handle up to 1M input tokens, the size of the passage collection is generally unbounded, making retrieval a necessity.. First, we retrieve a set of passages $D_{q}\subset D$ relevant to the query $q$ from the collection $D$. Then, we input the retrieved text and query to a language model $LM_{sum}$ to generate an abstractive summary $S=LM_{sum}(q,D_{q})$.
@@ 4 LightPAL: Proposed Method
We propose a retrieval method for addressing open-ended ODMDS queries. Our approach retrieves relevant and diverse passages while avoiding iterative retrieval and language model inference at runtime. As shown in Figure 1, The method consists of two phases: (1) Graph Construction: Pre-constructing a graph representing passage relationships using a large language model (LLM), linking passages with high conditional generation probabilities as potential context for a passage. (2) Retrieval: Retrieving an initial set of relevant passages with a conventional retriever, then employing random walk via Personalized PageRank (PPR) Jeh and Widom (2003a) on the pre-built graph. PPR calculates probabilities of reaching other passages from initial passages, retrieving highly probable ones as additional context without runtime language model inference.
Passage Graph Construction. The motivation behind our approach is that passages extracted from a larger document may not be self-explanatory and require context from other passages to be correctly understood. Even cohesive passages can benefit from related texts to deepen topic understanding. Naive retrieval systems that only consider query-passage relevance cannot capture these relationships. We aim to address this by creating a graph annotated with passage relationships using the conditional generation probabilities of LLMs. Concretely, for each ordered pair of passages $(d_{i},d_{j})$ in passage set $D$, we calculate the conditional generation probability of passage $d_{j}$ given passage $d_{i}$ as a context score using a language model:
@@
Passage Retrieval. Our proposed method consists of two phases for passage retrieval: 1) Initial Passage Retrieval: Retrieve passages $D_{init}$ directly relevant to the query using any sparse/dense retrieval model. 2) Context Passage Retrieval: Using the initial passages, additionally retrieve context passages $D_{context}$ from the passage graph $G$ that provide contextual information to the initial passages. Finally, we obtain $D_{q}$ as the union of $D_{init}$ and $D_{context}$. For context passage retrieval, we employ Personalized PageRank (PPR) Jeh and Widom (2003b). PPR calculates the probability of reaching other passages from the initial set $D_{init}$ via a random walk on the graph $G$, highly ranking passages cohesively referenced by many passages. Starting a random walk from the query-relevant initial passages, we expect to explore topics related to the query, and then retrieve highly informative context passages for those topics using PPR. This approach enables low-latency retrieval of relevant information without iterative search using LLM inference during runtime.
Story Meeting RAG-QA
# Queries 260 131 234
# Passages 1190 13074 3358
# PassLen. 3171 899 834
# SummLen. 1286 1081 1503
# Rel. 9.2 177.8 10.1
Rel. % 0.8% 1.4% 0.3%
Table 1: Statistics of the datasets. Abbreviations used: PassLen. (Average Passage Length in characters), SummLen. (Average Reference Summary Length in characters), Rel. (Average Relevant Passages), Rel. % (Relevant Passages Ratio).
@@ 5.1 Settings
Data. To evaluate our approach, we used three datasets: (1) Story dataset from ODSum Zhou et al. (2023), where passages are sections from independent stories, requiring retrieval of cohesive story sections relevant to the query. (2) Meeting dataset from ODSum, where queries are questions about meeting contents. We divided each long meeting transcript (approx. 8k tokens) into smaller chunks of around 1k characters, treating each chunk as a passage to avoid irrelevant information degrading summarization performance. If meeting transcript is labeled as relevant to query, chunks are also labeled as relevant. We used test set from both Story and Meeting. (3) RAG-QA dataset, derived from LFRQA Han et al. (2024), a benchmark for evaluating retrieval-augmented generation systems in long-form question answering. We linked LFRQA data with the LoTTE Santhanam et al. (2022) dataset to obtain relevant passage sets for each query. Queries are StackExchange question titles and Google search-autocomplete queries, while passages are StackExchange answers. We sampled up to 50 queries from five domains (lifestyle, recreation, science, technology, and writing), each with at least five related documents and answer lengths over 1k characters. We also added randomly sampled negative passages to create a comprehensive document collection. This dataset evaluates LightPAL’s performance in diverse, realistic information-seeking scenarios.
@@
Baselines for Passage Retrieval. To investigate the impact of retrieval methods on the final summary quality, we prepared two types of retrieval methods as naive baselines: a pre-trained dense retriever and a sparse retriever. The dense retriever embeds passages and queries into the same vector space and retrieves the top-K passages most similar to the query. We used the bge-large-en-v1.5  [^3] https://huggingface.co/BAAI/bge-large-en-v1.5 model for embedding as it is a top-tier model in Massive Text Embedding Benchmark  [^4] https://github.com/embeddings-benchmark/mteb. For the sparse retriever, we employed BM25 as an alternative baseline.
For Context Passage Retrieval, we compared LightPAL with PromptRank Khalifa et al. (2023), an unsupervised reranker for Multi-hop QA. PromptRank performs iterative reasoning and retrieval by traversing links between passages. It requires links between passages, so to enable a fair comparison, we employ the passage graph created by our method. For a fair comparison, both PromptRank and our proposed method initially used the top $|D_{init}|$ passages retrieved by either the dense retriever or BM25. Each method then retrieved the remaining $|D_{context}|=K-|D_{init}|$ context passages using their respective approaches. We set $|D_{init}|$ as top 60% of $K$.